User Tools

Site Tools


lbaops:lbamar2007:vt11c4holog

eVLBI data was being sent till about 084/1500

The log from experiment vt11c4 has been copied to: ftp.atnf.csiro.au:/pub/people/vlbi/incoming/

ANTAB and UVFLG files have been copied to the same place.

The setup for this experiment was the same as v190h

Agilent 4.2GHz

RX atten 10dB

LCP only

DAS profile: VSOP_HO.PRO

IFP #1 LCP 1634 - 1666 MHz

IFP #2 LCP 1666 - 1869 MHz

Channel 1, IFP #1, LO 1634-1650 MHz, LCP, Coherence = 163.9MHz

Channel 2, IFP #1, HI 1650-1666 MHz, LCP, Coherence = 165.5MHz

Channel 3, IFP #2, LO 1666-1682 MHz, LCP, Coherence = 167.1MHz

Channel 4, IFP #2-HI 1682-1698 MHz, LCP, Coherence = 168.7MHz

2007.084.22:25:19.70;“system is still coherent

2007.084.22:25:50.27;”no data being recorded.

2007.084.22:26:28.52;“observers were Warren Hankey and John Dickey.

2007.084.22:27:39.93;”system ran OK overnight.

Regards,

BRETT.

lbaops/lbamar2007/vt11c4holog.txt · Last modified: 2015/12/18 16:38 (external edit)